# HashAttention: Semantic Sparsity for Faster Inference

### v1

@@ 4.2 SCORE function for HashAttention
The bits are packed into an integer. We denote complete mapping function by $\phi_{\textrm{\small int \normalsize}}$
Implementation: The key signatures are computed and cached along with KV Cache. We store them in integer format. During the decoder hot path, we compute the query signature and perform bit operations to compute the hamming distance between queries and keys.
$\mathcal{H}(\phi_{kv}(\mathbf{k},\mathbf{v}),\phi_{q}(\mathbf{q}))$ (5)
@@
It should be noted that we can construct index on the obtained bit signatures for faster computation of SCORE + TOPK . We leave this for future exploration.
In our experiments, we focus on $\phi_{kv}$ that acts only on the key vector $\mathbf{k}$ and ignore the $\mathbf{v}$ vector or its features such as norm. As we will see in the next subsection, the importance of the token depends on the norm of its value vector $\mathbf{v}$. Incorporating $\mathbf{v}$ in $\phi_{kv}$ is easy and is left for future exploration.
Table 1: A head to head comparison of various sparsificiation techniques on datasets from LongBench from all the categories using the LLama-3.1-8B (128K context). InfLLM, DS and Quest use auxilliary memory for shorter representation of keys – noted in Aux-budget as bits per token. HashAttention while having smallest auxilliary budget outperforms all baselines on average.
Category $\rightarrow$ MQA SQA Summ Few-Shot Synthetic Code
@@
Full Model NA NA 54.97 55.17 25.20 91.65 99.50 55.04 63.59
Exact-Top NA 512 52.57 53.45 25.15 91.71 99.50 58.70 63.51
H2O NA 512 36.40 26.61 17.79 80.14 44.50 55.85 43.55
StreamLLM NA 512 31.53 27.98 17.85 50.16 12.50 58.15 33.03
InfLLM 256(pg=32,bit=16) 512 47.72 51.99 23.13 86.36 31.71 41.92 47.14
InfLLM 256(pg=16,bit=16) 512 48.67 53.09 22.76 86.85 33.54 42.97 47.98
DS 32(ch=16,bit=2) 512 50.99 50.57 23.42 90.61 99 57.41 62.00
DS 64(ch=16,bit=4) 512 52.86 53.44 23.75 89.88 99.00 56.36 62.55
DS 128(ch=8,bit=16) 512 40.86 43.21 21.74 87.06 84.00 52.99 54.98
Quest 32(pg=16,bit=2) 512 52.48 51.31 23.30 89.78 98.00 58.94 62.30
Quest 64(pg=16,bit=4) 512 52.71 54.35 24.15 91.82 98.50 59.90 63.57
Quest 128(pg=32,bit=16) 512 51.60 53.60 22.88 91.21 97.50 58.20 62.50
HashAttention 32 512 53.88 53.35 25.28 92.64 100.00 60.51 64.28
4.3 HashAttention module for adapting to pre-trained LLMs
@@
Consider $n$ KV embeddings $\mathcal{K},\mathcal{V}$ and query $\mathbf{q}$. Let $\mathbf{a}$ be the attention scores of tokens w.r.t query $q$. Then the contribution of a token $i$ towards the final output is proportional to
@@
The proof is presented in appendix D. Thus, the best set of $k$ tokens for attention computation are the ones that have the highest $a_{i}||\mathbf{v}_{i}||_{2}$. We use top tokens identified by this score to train our HashAttention modules for each of the attention.
We pose HashAttention training as a classification problem where each HashAttention has to predict the top-$k$ tokens of the attention head that it is associated with. We choose $k$ to be a small number such as $32$ or $64$. We use binary cross-entropy loss in a multi-class setting to train our functions $\phi_{kv}$ and $\phi$ with a standard Adam optimizer. Some important training details are mentioned below,
Class imbalance
As the context length increases, the class imbalance for classification increases. For instance at 64,000 context length, while using top-64 tokens to predict, only 0.1% of labels are of class $1$. We use class weights to resolve the issue of class imbalance. Since the imablance depends on the context length. We use the following formula to compute the class $1$ weights, parameterized with $\alpha$ and $\beta$.
@@
Soft-partitioning
We use the sign function to perform space-partitioning and obtain bits in $\phi$ mapping computation during inference as mentioned in the previous section. To train the functions, however, we use the tanh function in place of the sign function as a softer version of partitioning.
Training data
In order to train our HashAttention module, we use the openwebtext dataset  [^1] https://huggingface.co/datasets/Skylion007/openwebtext. We find that the choice of dataset does not play an important role in the quality of HashAttention. In fact, HashAttention trained on English language data performs well even on Chinese language tasks. We string together multiple examples from openwebtext to obtain longer sequences of required length for training. Additionally, we find that HashAttention trained on shorter sequences do not naturally scale to longer sequences. So, we need to train HashAttention with sequences of required length.
We run the LLM model on required length of tokens in inference mode in a chunk based fashion, i.e. the tokens are processed in chunks. At the end of each processing, all HashAttention modules are independently and locally trained for one step on queries from this chunk and previously cached KV embeddings.
@@ 5 Experiments
In this section, we evaluate HashAttention when adapted to existing LLM models. The section is organized as follows: first, we demonstrate the superiority of HashAttention in a head-to-head comparison against popular baselines. Next, we select the strongest baselines for a Pareto comparison (quality vs. token budget) of HashAttention and micro-benchmarking. Finally, we evaluate the efficiency of the HashAttention attention kernel.
Figure 4: Quality of models at different token budgets at an auxiliary budget of 32 bits/token.
Baselines:
@@
We use the following baselines: StreamingLLM (Xiao et al., 2023), H2O (Zhang et al., 2023), InfLLM (Xiao et al., 2024), DoubleSparsity (DS) (Yang et al., 2024), and Quest (Tang et al., 2024). We also include Exact-topk for quality comparison. For all sparsifying methods, we consistently include the first and most recent 128 tokens to ensure that the important parts of the prompt are always in context. The token budget refers to the number of heavy tokens retrieved.
InfLLM and Quest are page-based retrievers that represent each chunk using heuristic-based representative vectors to determine the chunk’s importance. These baselines have two parameters: page size (pg) and the number of representative vectors (rep). In InfLLM, we always set rep=4. While these baselines do not inherently incorporate further quantization, we apply additional quantization to reduce auxiliary memory usage and computational costs, as is done with DoubleSparsity (DS).
DS reduces the number of channels used to compute heavy tokens, with further quantization applied to these channels to minimize memory usage. DS has two parameters: the number of channels (ch) and the number of quantization bits (bit).
Figure 5: Recall (averaged across attention heads) of Layer 17 in Llama-3.1-8B model on an example from narrativeqa with context length of 64K tokens. Recall from other layers show similar patterns.
Evaluation Methodology
  • Prompt-offset: Previous works often apply full attention to prompts during the pre-fill phase and sparse attention during the decoding phase. However, this evaluation method is not ideal. The prompt typically consists of two parts: (1) the context (e.g., in passage retrieval datasets, the context is a collection of passages) and (2) the task (e.g., in passage retrieval datasets, the task includes a description of the passage to be retrieved). When full attention is applied to the entire prompt, the model processes both the task and the context with full attention. In practice, however, contexts are pre-filled and cached, while tasks must be processed later, where the quality of retrieval becomes critical. To simulate this scenario during evaluation, without relying on specific datasets, we apply sparse attention to the last 128 tokens of the prompt and all subsequent generations.
  • Auxiliary Memory: The $\mathrm{SPARSIFY}$ operation typically uses some metadata proportional to the number of tokens. For example, DS uses a label cache, QUEST uses $\mathrm{max}$ and $\mathrm{min}$ vectors for each chunk, and HashAttention uses bit signatures. Since memory is a significant bottleneck in deploying LLM models, we aim to evaluate different methods under the same auxiliary memory budget unless stated otherwise. To reduce the memory footprint of the InfLLM and QUEST baselines, we apply additional quantization to their representative vectors.
5.1 A head-to-head comparison of HashAttention against sparse-attention baselines
Models:
We use the latest Llama-3.1-8B model, which is trained for a 128K context length. It consists of $32$ decoder layers, each containing $32$ attention heads. All $1024$ attention heads are replaced with sparsifying modules.
Datasets: For this experiment, we randomly select one dataset from each category in the LongBenchmark (Bai et al., 2024b). Additionally, we use the first 200 samples from the repobench-p dataset.
The results are presented in the Table 1. We make the following observations:
  • We note that the exact-top model, which uses exact attention scores to determine top-k is competitive with the full model implying that predicting top attention scores is a viable direction for sparsifying attention.
  • Query agnostic sparsity-based methods such as H2O and StreamingLLM strictly underperform other approaches highlighting the importance of dynamic sparsity of contexts.
  • InfLLM uses dynamic sparsity but the heuristic of choosing representative vectors fails drastically in datasets such as passage retrieval.
  • DS and Quest are the strongest baselines for HashAttention. With a small enough page size for QUEST and a large enough number of channels in DS, they can compete with the full model. However, this comes at the cost of using larger auxiliary memory.
  • HashAttention outperforms all the baselines and matches full model quality while maintaining the least amount of auxiliary memory.
(a) Latency
(b) Latency ablation
Figure 6: Latency evaluation of HashAttention
5.2 Comparison of HashAttention across different token budgets
Settings:
We use the Llama-3.1-8B model for Pareto curve computation and select a range of datasets from LongBench. From Table 1, it is evident that Quest and DS are two strong baselines. In this Pareto curve analysis, we compare HashAttention against these baselines. For all methods, the auxiliary budget is set to 32 bits per token.
We make the following observations:
  • HashAttention outperforms other methods across various token budgets.
  • The quality gap at lower token budgets is particularly pronounced, with HashAttention demonstrating a significant advantage.
  • The quality achieved by the baselines at a given token budget can be matched by HashAttention at a much smaller token budget. This implies that the complexity of final sparse attention can be significantly reduced for HashAttention compared to other methods. As we will see, sparse attention is the major bottleneck in terms of latency, and reducing its complexity will result in substantial latency gains.
5.3 Recall of HashAttention
From the pareto curves, it is clear that Double Sparsity is the strongest baseline. We micro-benchmark Double Sparsity and HashAttention in their quality of retrieving top tokens at different context length. Figure 5 shows the mean recall of top-32 tokens at sparsity level 32$\times$ – meaning we only retrieve $1/32$ fraction of context-length. We note that the superior quality of HashAttention on benchmarks is a directly related to its superior retrieval quality.
5.4 Efficiency of HashAttention
As mentioned in the section 4, the ingenuity of different sparse attention methods lies in devising the SCORE function that lets us choose the sparsity pattern and perform reduced attention. The rest of implementation can be ported across different methods. In this section to evaluate latency of HashAttention, we use the sparse-fwd kernel from Double Sparsity and use pytorch based scoring function for HashAttention which is compiled using torch.compile. Figure 6(a) shows the relative latency of HashAttention compared to LightLLM(LightLLM, 2024) and FlashAttention from torch.scaled_dot_product_attention used in gpt-fast(GPTFast, 2024). FIgure 6(b) shows the split of time in different components of a sparse attention. We make the following observations,
  • The overall time of HashAttention is quite similar to that of Double Sparsity at the same sparsity. We observe 36×3-6$\times$3 - 6 × gains over LightLLM attention and 2.54.5×2.5-4.5$\times$2.5 - 4.5 × over flash attention from gpt-fast at a sparsity of 32$\times$.
  • The time overhead of computing SCORE is minimal as compared to the time spent in computing the actual sparse attention and performing TOPK .
  • As the number of tokens in sparse-fwd decreases (sparsity increases) the attention latency almost decreases linearly. This implies that there are direct gains from improving the recall of SCORE which lets us work with lesser number of tokens.
5.5 Discussion on latency
The sparse-fwd kernel from Double Sparsity does not implement sequence parallelism, as is the case with LightLLM attention and FlashAttention in gpt-fast. We are working on an implementation with sequence parallelism. With sequence parallelism, the sparse-fwd times may reduce making SCORE + TOPK the bottleneck.
6 Discussion
Our exposition of HashAttention is limited in several ways and requires further exploration for complete insights. Firstly, in the current experimental setup, while training the HashAttention modules, we targeted top-k tokens defined purely by attention scores and used only key vectors as input to the learned mapping functions. This was primarily done to compare all baselines with respect to their scoring recipes. We plan to continue exploring the inclusion of value vectors both as inputs to the mapping functions and as part of the target, which is easily achievable in HashAttention.
Additionally, further ablation studies are needed to explore various aspects of training HashAttention, including, but not limited to, training loss functions, targets for training, and so on. Currently, we retrieve a fixed number of tokens for all attention heads across all runs of the model. However, it may be valuable to retrieve different counts of tokens depending on the query. We plan to explore advanced retrieval mechanisms with HashAttention.
This paper focuses on adapting HashAttention to existing pretrained attention modules. However, if the existing pretrained full attention is inherently sparse and well approximated by HashAttention, it stands to reason that we can build an HashAttention that can be trained end-to-end to organically develop dynamic sparse attention as an integral part of the model, rather than as an afterthought. We plan to invest in developing end-to-end HashAttention next.

### v2

@@ 4.1 The SCORE function for HashAttention
The bits are packed into an integer. Let $\phi_{\textrm{\small int \normalsize}}$ be the compete mapping function.
Implementation: The integer signatures for key-values are computed and cached along with KV Cache. During the decoder hot path, we compute the query signature and perform bit operations to compute the hamming distance between the query and key-values.
$\mathcal{H}(\phi_{kv}(\mathbf{k},\mathbf{v}),\phi_{q}(\mathbf{q}))$
@@
In our experiments, we focus on $\phi_{kv}$ that acts only on the key vector $\mathbf{k}$ and ignore the $\mathbf{v}$ vector or its features such as norm. Incorporating $\mathbf{v}$ in $\phi_{kv}$ is left for future exploration.
Table 1: A head-to-head comparison of sparse-attention baselines on datasets from LongBench from all the categories using the LLama-3.1-8B-Instruct. InfLLM, DS and Quest use auxiliary memory for shorter representation of keys – noted in Aux-budget as bits per token. HashAttention while having smallest auxiliary budget outperforms all baselines on average. HashAttention is trained with mixture of openwebtext dataset and LongBench datasets (25 examples per dataset). Evaluation is done on first 175 examples (MFQA was excluded from training since it has only 150 samples)
Category $\rightarrow$ MQA SQA Summ FS-Learn Synthetic Code
@@
Full Model NA NA 54.83 55.17 24.91 91.31 100.00 55.07 63.55
Oracle(top) NA 512 52.10 53.45 25.14 91.39 100.00 58.49 63.43
H2O NA 512 36.62 26.61 17.85 80.75 43.43 55.55 43.47
StreamLLM NA 512 33.32 27.98 17.93 51.95 11.43 57.07 33.28
InfLLM 256(pg=32,bit=16) 512 47.75 51.99 23.14 88.32 33.00 42.99 47.87
InfLLM 256(pg=16,bit=16) 512 48.27 53.09 22.90 88.88 32.81 43.45 48.23
DS 32(ch=16,bit=2) 512 50.39 50.57 23.41 90.32 98.86 57.72 61.88
DS 64(ch=16,bit=4) 512 52.60 53.44 23.75 90.44 99.43 57.01 62.78
DS 128(ch=8,bit=16) 512 41.50 43.21 21.80 86.83 82.29 52.80 54.74
Quest 32(pg=16,bit=2) 512 53.13 51.31 23.01 90.15 98.29 58.29 62.36
Quest 64(pg=16,bit=4) 512 52.51 54.35 23.75 91.50 98.86 59.03 63.33
Quest 128(pg=32,bit=16) 512 51.60 53.60 22.94 90.53 97.71 57.06 62.24
HashAttention 32 512 54.08 53.35 25.08 92.41 100.00 59.98 64.15
[…]
Full 54.68 100 43.75 45.80 29.41 34.90 27.11 71.43 55.40 8.29 25.14 45.62 90.55 30.19 62.87 55.31 48.78 51.48
16x 53.65 100 42.82 43.52 29.16 34.99 26.68 70.29 51.92 1.85 24.75 42.97 92.14 27.93 64.89 60.49 48.00 51.08
Table 3: Llama3.1-8B-Instruct with HashAttention-16$\times$ on RULER@16K
niah_single_1 niah_single_2 niah_single_3 niah_multikey_1 niah_multikey_2 niah_multikey_3 niah_multivalue niah_multiquery vt cwe fwe qa_1 qa_2 Average
Full 100 100 100 100 100 100 99.25 98.75 99.2 50.5 90 87 54 90.66
16 $\times$ 100 100 100 99 90 96 98.5 98.75 93 62.4 87.33 82 57 89.53
4.2 HashAttention for adapting to pre-trained LLMs
@@
Consider $n$ KV embeddings $\mathcal{K},\mathcal{V}$ and query $\mathbf{q}$. Let $\mathbf{a}$ be the attention scores of tokens w.r.t query $q$. Then the contribution of a token $i$, assuming no other token is used, towards the final output is proportional to
@@
The proof is presented in appendix B.1. We can use top tokens identified by this score to train our HashAttention modules for each attention. We pose HashAttention training as a classification problem where each embeddings generated in HashAttention are used to predict the top-$k$ tokens of the attention head that it is associated with. We choose $k$ to be a small number such as $64$. We use binary cross-entropy loss in a multi-class setting to train our functions $\phi_{kv}$ and $\phi_{q}$ with a standard Adam optimizer. We mention some important details below,
Class imbalance As the context length increases, the class imbalance for classification increases. For instance at 64,000 context length, while using top-64 tokens to predict, only 0.1% of labels are of class $1$. We use class weights to resolve the issue of class imbalance. Since the imablance depends on the context length. We use the following formula to compute the class $1$ weights, parameterized with $\alpha$ and $\beta$:
@@
Soft-partitioning While we use the sign function to perform space-partitioning and obtain bits in $\phi$ mapping computation during inference, we use the tanh function in place of the sign function as a softer version of partitioning while training.
Training dataset We find that HashAttention trained on a generic dataset works well on a variety of tasks. To further improve the quality, we can finetune HashAttention on task-specific data. Additionally, we find that HashAttention trained on shorter sequences (say $<=64K$) do not naturally scale to longer sequences (e.g. $128K$). So, HashAttention needs to be trained for appropriate context-lengths.
We run the LLM model in inference mode in a chunk-based fashion, i.e. the tokens are processed in chunks. At the end of each inference, all HashAttention modules are independently and locally trained for one step on queries from this chunk and previously cached KV embeddings.
4.3 Theoretical underpinnings of HashAttention
In this section, we briefly explain why HashAttention works. Lemma 4.1 shows that analyzing token importance in isolation, the contribution of $i^{th}$ token is proportional to $a_{i}||\mathbf{v}_{i}||_{2}$ where $a$ and $\mathbf{v}$ are attention score and value vector respectively. Identifying tokens with this true score can be cast as a maximum inner product search problem as shown below,
Lemma 4.2.
Consider a KV Cache $\mathcal{K},\mathcal{V}$, and query vector $\mathbf{q}$. The ordering imposed on scores $a_{i}||\mathbf{v}_{i}||_{2}$ for $i^{th}$ token is same as the ordering imposed by inner product $\langle[\mathbf{q},1],[\mathbf{k},log(||\mathbf{v}||_{2})]\rangle$ where [] denotes concatenation.
The proof is presented in appendix B.1. The Maximum inner product search problem can be reduced to a cosine-similarity search problem (Shrivastava & Li, 2014; Neyshabur & Srebro, 2015) using asymmetric transformation on query and key-values. Combining Lemma 4.1 and Lemma 4.2 along with results from (Neyshabur & Srebro, 2015), we can equivalently determine ordering based on cosine similarity between the following vectors,
$\psi_{kv}(\mathbf{k},\mathbf{v})=\left[\mathbf{k},\log(||\mathbf{v}||_{2}),\sqrt{M^{2}-||[\mathbf{k},log(||\mathbf{v}||_{2})]||^{2}}\right]$
where M is the maximum norm of all vectors $[\mathbf{k},\log(||\mathbf{v}||)$
$\psi_{q}(\mathbf{q})=\left[\mathbf{q},1,1\right]$
Cosine similarity can be approximated using bit signatures obtained by random signed projections (Gionis et al., 1999). HashAttention is motivated by this theoretical understanding of similarity search. Furthermore, we use learnable functions in place of $\psi$ to exploit the patterns in query and key distribution and obtain succinct bit signatures. Using independent encoding functions for key-values and queries helps in solving the MIPS problem.
4.4 Efficiency of token level sparsity.
As opposed to methods such as Quest or InfLLM that use block sparsity in attention, HashAttention uses token-level sparsity. We want to note that token-level sparsity is unavoidable when trying to implement sparse attention during inference time when training assumes full attention, because important tokens, defined by higher attention scores, may not appear in the same block, and not using some important tokens to adhere to an imposed block size will only lead to an inferior top-k approximation. Interestingly, and contrary to popular belief, token-level sparsity does not lead to system inefficiency.
HashAttention implementation is built upon the vLLM page attention framework (Kwon et al., 2023), where each token corresponds exactly to one page (i.e., page size = 1). After identifying top-k tokens, we use their indices without physically moving or gathering these tokens into a separate memory buffer. The page attention kernel explicitly utilizes these indices to selectively compute attention only for the specified tokens, efficiently ignoring irrelevant tokens. The page attention kernel is highly optimized for GPU memory access patterns. Due to the GPU cache line size of 128 bytes, optimal memory bandwidth utilization is achieved as long as contiguous data access meets or exceeds this cacheline size. It is important to note that many state-of-the art inference framework implement attention using paged-attention backbone with page size $1$. They find that with correct optimization, the efficiency does not depend on page size  (Ye et al., 2024).
@@ 5 Experiments
In the experiments below, we use two versions of HashAttention. HashAttention is trained using openwebtext(Gokaslan et al., 2019) dataset with samples concatenated together to obtain the required context length. HashAttention is further finetuned from HashAttention for specific dataset.
@@
baselines: StreamingLLM (Xiao et al., 2023), H2O (Zhang et al., 2023), InfLLM (Xiao et al., 2024), DoubleSparsity (DS) (Yang et al., 2024), and Quest (Tang et al., 2024). We also include Oracle which predicts exact top-k. For all methods, we include the first and most recent 128 tokens to ensure that the important parts of the prompt are always in context. In cases where methods use auxiliary memory (e.g. label cache in DS and bit signatures in HashAttention), we denote it as bits per token per head(PTPA). The token budget refers to the number of heavy tokens retrieved. InfLLM and Quest are page-based retrievers. These baselines have two parameters: page size (pg) and the number of representative vectors (rep). In InfLLM, we always set rep=4. While these baselines do not inherently incorporate further quantization, we apply additional quantization to reduce auxiliary memory usage, as is done with DoubleSparsity (DS). DS has two parameters: the number of channels (ch) and the number of quantization bits (bit). Baselines are further explained in Appendix A.
prompt-offset: To simulate a scenario of preprocessed long context with relatively smaller prompts, without relying on specific datasets, we apply sparse attention only to the last offset tokens of the complete prompt and all subsequent generations. We use offset of 128 for Longbench and 32 in RULER.
Auxiliary Memory: Since memory is a significant bottleneck in deploying LLM models, we aim to evaluate different methods under the same auxiliary memory budget of 32 bits PTPA unless stated otherwise.
Models and datasets: We use Llama-3.18B-Instruct(Dubey et al., 2024) and Mistral-7B-Instruct-v0.3(Jiang et al., 2023) models and Longbench(Bai et al., 2024b) and RULER(Hsieh et al., 2024) datasets in our evaluation. All attention heads are replaced with sparse attention.
Efficiency frameworks: We use GPT-FAST(GPTFast, 2024) and FlashDecode(Dao, 2023) to test improvements offered by HashAttention.
5.1 Quality with HashAttention
We present the following experiments to evaluate the quality of HashAttention.
A head-to-head comparison of HashAttention at fixed budget For this experiment, we randomly select one dataset from each category in the LongBench and use the first 175 samples (the last 25 samples are added to the training set). We set a budget of 512 pivotal tokens for all baselines. The results are presented in the Table 3.
Pareto curves To get a broader comparative picture, we use four datasets (Passage Retrieval (English), TriviaQA, HotpotQA, and MultifieldQA (English) ) from Longbench and six datasets (niah, vt, cwe, fwe, qa1, and qa2) from RULER (16K) to perform a more elaborate Pareto analysis. We compare against Quest and DS at 32 PTPA auxiliary budget. Unless explicitly mentioned, we use HashAttention (and not HashAttention ). These results are presented in Figure LABEL:fig:pareto.
Full benchmark at 16$\times$ sparsity We also present HashAttention at 16×16$\times$16 × sparsity on full LongBench and RULER benchmarks for Llama-3.1-8B-Instruct. The results are presented in Table 3 and Table 3.
Micro benchmarking the recall rates To show that the improvements in quality come from the improvement in pivotal token retrieval, we show the recall rates of top-32 tokens at different context lengths for the Llama-3.1-8B-Instruct model. The result is presented in Figure LABEL:fig:recall in the appendix.
We make the following observations,
  • Table 3 presents the landscape of different sparse-attentions. H2O and StreamingLLM fail due to the incorrect fixed sparsity assumption. InfLLM uses dynamic sparsity, but the heuristic-based scoring does not generalize well over all tasks leading to poor performance. Quest and DS perform relatively well on all tasks. However, it needs more auxiliary memory to obtain good results. Overall, HashAttention outperforms all the baselines while using minimum auxiliary memory.
  • Figure LABEL:fig:pareto, Table 3, and Table 3 show that HashAttention can significantly reduce the number of tokens in attention computation. HashAttention is almost 4×4$\times$4 × more token-efficient compared to Quest and DS at an auxiliary budget of 32 bits/token. Overall, HashAttention seems to preserve most of the model quality at 16$\times$ sparsity (0.78 points less on Longbench, 0.4 points without the worst performing dataset, and 1.13 points less on RULER). The results can be further improved using HashAttention.
  • Certain levels of sparsity seem to improve quality over the full model in certain tasks. This could be potentially explained as sparse attention avoids attention to distractor tokens.
  • HashAttention trained on English data also performs reasonably well on Chinese datasets in LongBench, although with a higher drop in quality. (2.66 points on average. See table 5 in the appendix.) The results on coding tasks (lcc and repobench-p in Longbench) are especially better than the full model. Coding tasks might be more sensitive to distractors than general English tasks.
  • The recall of pivotal tokens shown in Figure LABEL:fig:recall (in appendix) gives an insight into why HashAttention outperforms other baselines. The reason is that HashAttention is better at retrieving pivotal tokens.
5.2 Efficiency with HashAttention
To evaluate the efficiency that HashAttention has to offer, we perform the following experiments. In these experiments, we use 3-layered MLP (128x128-128x128-128x32) for mappings learned in HashAttention. Using smaller MLPs will improve latency results for HashAttention. We use 32×32$\times$32 × sparsity for these experiments.
  • We incorporate HashAttention in the GPT-FAST attention kernel and FlashDecode attention kernel and measure the latency improvements.
  • We incorporate HashAttention in the GPT-FAST inference system to measure the end-to-end improvement in throughput.
We make the following observations.
  • The matrix multiplication in mapping functions used to obtain bit-signatures dominates the latency up to 8K context length in GPT-FAST and up to 65K in Flash Decode. Since the main difference between the two attention implementations is that FlashDecode uses sequence parallelism, whereas GPT-FAST does not, this behavior is expected. Post 4K context length for GPT-FAST and 65K context length for FlashDecode, HashAttention improves the latency for the attention. We can see up to 4.3$\times$ improvement in GPT-FAST and 2.54$\times$ improvement in FlashDecode.
  • The throughput improvements can be seen starting at 8K context length. The throughput improvement ( ratio of the throughputs) starts declining as sequence length increases. This can be attributed to TOPK operation becoming expensive at higher sequence lengths. In this experiment, we see the best throughput gains of 3.12$\times$ occur at a sequence length of 32K for 32$\times$ sparsity.
5.3 Ablations
In this section, we present ablations. The tables / figures for this section are presented in Appendix.
HashAttention Bit Signatures vs. LSH Bit Signatures:
We compare LSH bit signatures (Gionis et al., 1999), obtained via random signed projections, with learned bit signatures used in HashAttention. We find that LSH requires significantly more bits to perform reasonably well on top-k prediction. This inefficiency stems from the agnostic nature of LSH. The results are presented in Table 8. While HashAttention performs well with 32-bit signatures, LSH fails to achieve comparable performance even with more than 1000 bits.
Note that this is not a direct comparison to MagicPig (Chen et al., 2024), as MagicPig builds LSH tables from bit signatures for sparse retrieval and modifies the attention computation to incorporate the sampling view of LSH. Nevertheless, it provides a reasonable indication of the auxiliary memory used by MagicPig to achieve sparse attention.
Query and Key distributions under HashAttention embedding
While HashAttention does not directly aim to resolve the out-of-distribution issue between queries and keys, the learned mappings bring the distributions closer due to the nature of the training loss. We measure the empirical cosine similarity between each query and its top-32 tokens using the original embeddings, HashAttention (signed) embeddings, and HashAttention (tanh) embeddings used during training. We find that the cosine similarity between key and query embeddings is significantly higher when using HashAttention embeddings compared to the original embeddings. The results are presented in Table 7.
Quality vs. bit-width in HashAttention
As expected, we find that using more bits helps with the quality of top-k prediction. The cross-entropy loss of top-32 prediction using different bit-width is presented in Table 9.
Latency of SCORE computation vs. full inner product.
Since increasing bit-widths can improve attention quality, it is reasonable to ask how many bits can be used before the computational advantage of HashAttention disappears. We present this ablation in Table 6. We find that even for bit-widths as large as 512, the latency of SCORE computation using our optimized kernel—which employs bitwise operations—is still lower than that of full inner product computation for 128-dimensional vectors with float16 precision.
6 Limitations and Discussion
We discuss the limitations of the proposal and evaluation here. HashAttention needs to be trained at least once. As opposed to training-free methods this is a limitation since every model needs its own HashAttention. This is, however, not very different from the calibration used in DS. To obtain better results, we can use further finetuning on the task-specific data. Practically, this is okay for users who care about a specific application of LLM. In terms of evaluation, we restrict our evaluation to situations where KV Cache is on the GPU. For extremely long context situations, KV Cache needs to be stored on the CPU RAM, and evaluation of HashAttention in such a scenario using flash-decode would be most valuable. However, such an evaluation is out of the scope of this paper. Having said that, HashAttention clearly shows superior quality, larger sparsity, and lower auxiliary memory compared to different top-k-based sparse attentions under identical conditions.